# Scene-Text Grounding for Text-Based Video Question Answering

### v1

@@ III-C2 Optimization
where $\mathrm{sim}(\cdot)$ denotes cosine similarity function and temperature coefficient $\tau$ = 0.1.
@@
IV Dataset
IV-A ViTXT-GQA  Construction
Dataset Filtering. We only label the validation and test sets for a weakly-supervised setting. Before the annotation, the questions whose answers do not involve scene texts or involve multiple text regions on a single frame are removed. Moreover, the questions whose answer texts are blurry or occluded in the video frame are also removed. Consequently, 343/386 videos and 981/1,074 questions on val/test sets are to be annotated.
@@ IV-C Dataset Comparison
Comparison with TextVideoQA. NewsVideoQA [17], RoadTextVQA [31], and M4-ViteVQA [42] aim to predict textual answers. ViTXT-GQA  differs in two main aspects: 1) it provides visual evidence to support the answers, and 2) it extends the VQA setting by allowing visual answers. This satisfies more practical applications and facilitates better diagnostics of model performance. For example, are the incorrect answer predictions due to scene text recognition errors or overly strict answer evaluation? ViTXT-GQA is also more challenging because: 1) models need to achieve multiple objectives (i.e., grounding and QA) while maintaining consistency, and 2) models must accomplish not only temporal grounding but also spatial grounding, which is challenging to locate scene text that answers the question in visually rich videos.
@@
V Experiments
$\bullet$ Q1: How effective is our T2S-QA model?
$\bullet$ Q2: What is the contribution of each component in T2S-QA?
$\bullet$ Q3: How does scene-text recognition affect TextVideoQA?
TABLE III: Performance comparison on ViTXT-GQA test set. Models use ABINet [11] by default for scene text recognition. We additionally upgrade to CLIPOCR [34]. The best and second-best results are boded and underlined respectively.
@@ V-A Settings
Implementation Details. We adopt a pre-trained BERT-Base [7] model to encode the question. The encoder is also finetuned. For video embedding, each frame is encoded by a ViT-L [9] model pre-trained on ImageNet-21k. We standardize the video frame rate to 10 frames per second and set frame number $T$ = 64, OCR token number $S$ = 15 in each frame, the maximal question length $L$ = 20, and the common hidden dimension $d$ = 768. For answer decoder and model optimization, we set the step number of answer decoding to $L_{a}$ = 12 and the trade-off hyperparameter $\lambda$ = 100. We collect the top 5,000 frequent words from the answers in the training set as a fixed vocabulary. During training, we set batch size to 48 and train for a maximum of 24,000 iterations. Our model is trained using the Adam optimizer, with a learning rate of 1e-4 and a staircase learning rate schedule, where we multiply the learning rate by 0.1 at 10,000 and at 20,000 iterations.
TABLE IV: Ablation study of temporal-spatial grounding.
@@
TABLE V: Ablation study of the input of answer decoder.
@@
Evaluation. We consider QA and answer grounding evaluation. For QA, we adopt Accuracy (Acc) in percentage $\in[0,100]$ and Average Normalized Levenshtein Similarity (ANLS) [2] $\in[0,1]$. Acc considers a predicted answer to be correct if and only if it matches the ground-truth answer word by word. ANLS measures the character-level compositional similarity between the predicted answer and the ground truth. For answer grounding, we adopt Intersection over Union (IoU) between the grounded and the annotated OCR token boxes. We consider IoU at two thresholds of 0.3 and 0.5 for better analysis. To evaluate the joint answer grounding and QA, we consider grounded QA accuracy (GQA). Specifically, we report GQA@0.5 and GQA@0.3 metrics which stand for the percentages of correctly answered questions with grounding IoU $\geq$ 0.5 and 0.3 respectively. Finally, we set two evaluations considering the severe challenge of video scene-text grounding. The first is a normal evaluation of the Top 1$\times$1, which measures the Top-1 grounded scene text region within the Top-1 temporally localized video frame. The second is Top 5$\times$5, which relaxes the requirements from Top-1 to Top-5 for both spatial and temporal grounding, i.e., measuring if the identified Top-5 frames and the Top-5 regions within each frame contains the ground-truth scene-text region.
V-B Main Comparison (RQ1)
As there is no existing method for Grounded TextVideoQA. We adopt several approaches from TextVQA, TextVideoQA, and VideoQA as baselines. For a fair comparison, all methods share feature encoding and answer decoding backbones.
M4C [16] is an image-based TextVQA model. We adapt it to the video domain by randomly sampling a single frame. In practice, we sample the first, middle, and end frames, and average their results for reference. To obtain spatial grounding results, we employ post-hoc attention analysis as described in Eqn. 4 $\sim$ 5. We select Top $K_{2}$ OCR tokens according to the attention distribution of Eqn. 5. M4C utilizes the sampled frame features and OCR token features for the QA prediction.
The main results in Tab. III show that our model T2S-QA consistently outperforms all baselines under different evaluation metrics and settings. Such strength is unique to T2S-QA, as no single baseline method outperforms others across all metrics for both scene-text grounding and TextVideoQA. More specifically, T2S-QA wins more under the Top-5 setting versus the Top-1, especially for grounding. By looking at performances of all other methods under the Top-1 setting, we believe that it is very challenging to precisely localize the exact scene-text region for a particular answer, which leads to a relatively marginal performance gap among models.
@@ V-B Main Comparison (RQ1)
TABLE VI: Performance with different numbers of video frames. Box Cov. (%) is the average coverage percentage of annotated boxes w.r.t the sampled video frames.
@@ V-C Ablation Study (RQ2)
Effectiveness of Temporal-to-Spatial Grounding. In Tab. IV, we respectively study the effectiveness of T2S-QA by removing temporal and spatial grounding (“w/o TSG”), only removing temporal grounding (“w/o TG”), and only removing spatial grounding (“w/o SG”). Specifically, “w/o TG” variant uses the grounded OCR tokens in all video frames as positive OCR tokens and the others as negative OCR tokens; “w/o SG” variant uses all OCR tokens of the positive video frame as the positive OCR token, and all OCR tokens of the negative video frame as the negative OCR token; “w/o TSG” variant represents reserving all video frames and OCR tokens for answering, i.e., without answer grounding.
Tab. IV shows that both spatial and temporal grounding are key for better QA performance. We note that the model variant “w/o SG” performs the worst in both answer grounding and grounded QA. This suggests that only learning the differences between video frames is insufficient to capture subtle local changes in scene text within a frame, making it challenging to achieve accurate answer grounding. Moreover, compared with T2S-QA, “w/o TG” also shows a significant decrease of 13.29% on IoU@0.5 and a 1% decrease in GQA@0.5. This indicates that directly performing one-stage grounding of scene text across the entire video yields suboptimal results.
Study Inputs of Answer Decoder. Tab. V shows that when removing OCR tokens (“w/o O”) both T2S-QA’s grounding and QA results degenerate significantly by 11.16% on IoU@0.5 and 4.76% in Acc respectively. This demonstrates the significance of OCR token grounding for answer prediction. Additionally, we show that both the frame appearance F feature and question Q help the performance.
@@
Different Numbers of Video Frames. We conduct experiments with four different numbers of sampled video frames, i.e., $T$ = 32, 48, 64, and 80. Tab. VI shows that uniformly sampling 80 frames can cover over 96% of the annotated scene-text regions. However, the best results are achieved at frame number of 64. We speculate that learning with too many frames make it harder to identify the very limited positive frames that carry the needed scene texts.
TABLE VII: Study of scene-text grounding and recognition in Grounded TextVideoQA on ViTXT-GQA test set.
@@
Figure 6: Case-Study of T2S-QA on ViTXT-GQA test set. The green and yellow boxes in video frames denote the ground-truth boxes and the model grounded OCR tokens. The blue bar denotes the annotated temporal segment.
@@ V-D In-Depth Analysis (RQ3)
Our previous experiments show that models obtain low QA accuracy even with relatively higher grounding behavior. We speculate that the reasons are two folds: poor scene-text recognition and ineffective answer evaluation. The former hints that the OCR system fails to convert the grounded scene-text into textual answers (e.g. M vs. M.P.H). The latter indicates that the predicted textual answer does not exactly match the ground-truth one (e.g. 30 vs. 30 M.P.H in Fig. 1).
Effect of Scene Text Recognition. Scene text detection and recognition are two essential components of OCR systems. In this work, we analyze scene-text recognition as it is responsible for converting the grounded scene text into textual answers. To study the effect of scene-text recognition, we consider both 1) the variation of scene-text recognition systems and 2) using ground-truth scene-text as inputs of the answer decoder. For the former, we upgrade the default scene-text recognition method ABINet to the current SOTA method CLIPOCR [34]. As shown in Tab. VII (Bottom), this indeed consistently boosts the model performance, indicating that better scene-text recognition brings better TextVideoQA results. For the latter, we isolate the effect of incorrect grounding by directly inputting the ground-truth scene text of each QA pair to the answer decoder. Surprisingly, we only observe marginal improvement of accuracy (1.76%) even with perfect grounding results. The experiments thus demonstrate the severe limitation of existing methods for scene text recognition in videos.
@@
Human Performance and Upper Bound. To discern the error from ineffective answer evaluation, we sample around 20% test data which contains 200 questions and 163 videos and conduct a human study by asking volunteers to answer the questions and ground the related scene text areas. Table VII (Top) shows that even human (we assume that humans are with perfect scene-text recognition.) can only achieve an accuracy of 64.32%. A further investigation of the the wrong answers suggests that an absolute 25% out of 36% of they are acceptable and should be regarded as correct. For example, “35131.86” vs. “35,131.86”. The results thus suggest that stringent string matching is ineffective in assessing TextVQA models. Yet, human’s much higher grounding performance suggests that grounding annotation has much less subjectivity and demonstrates our significance to directly evaluate the scene-text regions. Additionally, the large gap between our model and human performance calls for more research towards trustworthy TextVideoQA.
@@
We also report an upper-bound result under the current scene text detection system [35] and recognition system [34]. The result is obtained by a brute-force search over all the detected scene text regions and recognized texts. The results in Table VII (Top) suggest that under the current OCR systems, even a brute-force search can only achieve a grounding accuracy of 56.52%. Note that T2S-QA achieves 28.60%. This reflects the effectiveness of our grounding approach, though the absolute accuracy is not so high. Additionally, we find that the upper-bound is inferior to human in grounding but exceeds human in QA accuracy. This reveals the inadequacy of the existing OCR system in scene-text detection and the challenge of effectively evaluating textual answers because of human subjectivity.

### v2

@@ III-C2 Model Optimization
where $\mathrm{sim}(\cdot)$ denotes cosine similarity function, and we empirically set temperature coefficient $\tau$ = 0.1 following [6, 26].
@@
IV Dataset
IV-A ViTXT-GQA  Construction
Dataset Filtering. We only label the validation and test sets for a weakly-supervised setting. To ensure the clarity of answer grounding, the questions whose answers do not involve scene texts or involve multiple text regions on a single frame are removed before the annotation. Moreover, the questions whose answer texts are blurry or occluded in the video frame are also removed. Consequently, 343/386 videos and 981/1,074 questions on val/test sets are to be annotated.
@@ IV-C Dataset Comparison
Comparison with TextVideoQA. NewsVideoQA [21], RoadTextVQA [41], and M4-ViteVQA [55] aim to predict textual answers. ViTXT-GQA  differs in two main aspects: (1) it provides visual evidence to support the answers, and (2) it extends the VQA setting by allowing visual answers. This satisfies more practical applications and facilitates better diagnostics of model performance. For example, are the incorrect answer predictions due to scene text recognition errors or overly strict answer evaluation? ViTXT-GQA is also more challenging because: (1) models need to achieve multiple objectives (i.e., grounding and QA) while maintaining consistency, and (2) models must accomplish not only temporal grounding but also spatial grounding, which is challenging to locate scene text that answers the question in visually rich videos.
@@
TABLE III: Performance comparison on ViTXT-GQA test set. Models use ABINet [12] by default for scene text recognition. We additionally upgrade to CLIPOCR [45]. The best and second-best results are bolded and underlined respectively.
@@
TABLE IV: Performance of MLLMs (zero-shot) on ViTXT-GQA. The evaluation is conducted under the Top 1$\times$1 setting.
  Method IoU@0.5 Acc ANLS GQA@0.5
          Question Answering
GPT-4o-mini w/ Standard - 34.19 0.5070 -
Qwen2-VL w/ Standard - 41.34 0.5649 -
Question Answering with CoT Prompt
GPT-4o-mini w/ CoT - 37.99 0.5574 -
Qwen2-VL w/ CoT - 41.90 0.5663
Question Answering with Grounded CoT Prompt
GPT-4o-mini w/ Grounded CoT 0.19 34.46 0.5058 0.00
Qwen2-VL w/ Grounded CoT 2.14 26.07 0.3628 1.02
Qwen2-VL w/ One-shot Grounded CoT 2.94 26.98 0.3664 1.32
T2S-QA (CLIPOCR) 3.21 9.87 0.1477 0.88
Figure 5: Prompts for MLLMs to perform ViTXT-GQA.
V Experiments
$\bullet$ Q1: How effective is our T2S-QA model?
$\bullet$ Q2: What is the contribution of each component in T2S-QA?
$\bullet$ Q3: How does scene-text recognition affect TextVideoQA?
@@ V-A Settings
Implementation Details. We adopt a pre-trained BERT-Base [8] model to encode the question. The encoder is also finetuned. For video embedding, each frame is encoded by a ViT-L [10] model pre-trained on ImageNet-21k. We standardize the video frame rate to 10 frames per second and set frame number $T$ = 64, OCR token number $S$ = 15 in each frame, the maximal question length $L$ = 20, and the common hidden dimension $d$ = 768. For answer decoder and model optimization, we set the step number of answer decoding to $L_{a}$ = 12 and the trade-off hyperparameter $\lambda$ = 100. We collect the top 5,000 frequent words from the answers in the training set as a fixed vocabulary. During training, we set batch size to 48 and train for a maximum of 24,000 iterations. Our model is trained using the Adam optimizer, with a learning rate of 1e-4 and a staircase learning rate schedule, where we multiply the learning rate by 0.1 at 10,000 and at 20,000 iterations. The best checkpoint is selected using the validation set accuracy.
Evaluation. We consider QA and answer grounding evaluation. For QA, we adopt Accuracy (Acc) in percentage $\in[0,100]$ and Average Normalized Levenshtein Similarity (ANLS) [2] $\in[0,1]$. Acc considers a predicted answer to be correct if and only if it matches the ground-truth answer word by word. ANLS measures the character-level compositional similarity between the predicted answer and the ground truth. For answer grounding, we adopt Intersection over Union (IoU) between the grounded and the annotated OCR token boxes. We consider IoU at two thresholds of 0.3 and 0.5 for better analysis. To evaluate joint QA and answer grounding, we consider grounded QA accuracy (GQA) [49]. Specifically, we report GQA@0.5 and GQA@0.3 metrics which stand for the percentages of correctly answered questions with grounding IoU $\geq$ 0.5 and 0.3 respectively. Finally, we set two evaluations considering the severe challenge of video scene-text grounding. The first is a normal evaluation of the Top 1$\times$1, which measures the Top-1 grounded scene text region within the Top-1 temporally localized video frame. The second is Top 5$\times$5, which relaxes the requirements from Top-1 to Top-5 for both spatial and temporal grounding, i.e., measuring if the identified Top-5 frames and the Top-5 regions within each frame contains the ground-truth scene-text region.
V-B Main Comparison (RQ1)
As there is no existing method for Grounded TextVideoQA. We adapt several approaches from TextVQA, TextVideoQA, and VideoQA as baselines in Tab. III. For a fair comparison, all methods in Tab. III share feature encoding and answer decoding backbones. Besides, we evaluate the zero-shot performance of the current advanced MLLMs Qwen2-VL [44] and GPT-4o-mini [34] in Tab. IV. We elaborate on these model details as follows.
M4C [20] is an image-based TextVQA model. We adapt it to the video domain by randomly sampling a single frame. In practice, we sample the first, middle, and end frames, and average their results for reference. To obtain spatial grounding results, we employ post-hoc attention analysis as described in Eqn. 7 $\sim$ 8. We select Top $K_{2}$ OCR tokens according to the attention distribution of Eqn. 8. M4C utilizes the sampled frame features and OCR token features for the QA prediction.
Qwen2-VL [44] is a cutting-edge open-source MLLM that supports video inputs, excelling in VQA and visual grounding tasks. By integrating temporal and spatial information, it delivers strong performance in complex spatio-temporal reasoning and grounding scenarios. To explore the performance of Qwen2-VL on the Grounded TextVideoQA task, we design a dedicated Chain-of-Thought prompt (CoT) (Fig. 5) to explore its QA and grounding capabilities. The CoT prompt guides the model in identifying the most relevant video frame and scene text bounding box for answering the question.
GPT-4o-mini [34] is a closed-source, lightweight variant of the GPT-4 family, designed for efficiency while maintaining robust multimodal reasoning capabilities. Although it performs well in VQA, its visual grounding ability is constrained by weaker spatio-temporal reasoning. As with Qwen2-VL, using the designed CoT prompt in Fig. 5, we also evaluate GPT-4o-mini’s performance on the Grounded TextVideoQA task.
TABLE V: Ablation study of temporal-spatial grounding.
@@ V-B Main Comparison (RQ1)
TABLE VI: Ablation study of the input of answer decoder.
@@
The main results in Tab. III show that our model T2S-QA consistently outperforms all baselines under different evaluation metrics and settings. Such strength is unique to T2S-QAcompared with other baseline methods. More specifically, T2S-QA wins more under the Top-5 setting versus the Top-1, especially for grounding. Examining the performance of all other methods under the Top-1 setting, we find that accurately localizing the precise scene-text region corresponding to a specific answer is highly challenging. This also explains the relatively marginal performance gap among models.
In Tab. IV, we evaluate the performance of MLLMs on ViTXT-GQA through three setups: direct QA (“w/ Standard”) in Fig. 5 (a), QA with Chain-of-Thought (CoT) prompt (“w/ CoT”) in Fig. 5 (b), and scene-text grounded QA with Chain-of-Thought prompt (“w/ Grounded CoT”) in Fig. 5 (c). Our findings are as follows: (1) MLLMs face severe challenges in scene-text spatio-temporal grounding despite their good QA performance. The results show that under the “Grounded CoT” strategy, GPT-4o-mini [34] and Qwen2-VL [44] achieve low IoU@0.5 scores of 0.19% and 2.14%, respectively, failing behind T2S-QA. (2) CoT prompting improves QA accuracy. Switching from “w/ Standard” to “w/ CoT”, GPT-4o-mini [34] shows a 3.80% increase in Acc, demonstrating CoT’s effectiveness in enhancing reasoning. To further explore the CoT prompt, we adopt a representative QA pair as the one-shot CoT prompt in Fig. 5 (c). The results demonstrate a modest improvement in both textual answer accuracy and visual grounding performance. (3) The performance of explicit grounded reasoning and implicit grounded reasoning is inconsistent. When transitioning from “w/ Standard” to “w/ Grounded CoT”, Qwen2-VL [44] suffers a 15.27% decrease in Acc, indicating the inability to effectively integrate grounding in the QA process. We further identify key QA failure cases, including reliance on irrelevant scene text or hallucinated visual content and difficulty recognizing occluded or partially visible text, underscoring the challenges and future directions in scene text-aware video QA.
TABLE VII: Ablation study of different scene text features. Linguistic features include FastText feature and PHOC feature. Spatial features include bounding box features, Temporal features include temporal id features and track id features.
  Linguistic Spatial Temporal IoU@0.5 Acc ANLS GQA@0.5
  - - 24.64 10.30 0.1414 3.05
- 28.31 10.39 0.1471 2.60
28.60 10.45 0.1520 3.50
TABLE VIII: Performance with different numbers of video frames. Box Cov. (%) is the average coverage percentage of annotated boxes w.r.t the sampled video frames.
@@
TABLE IX: Performance with different numbers of transformer layers.
  Variant IoU@0.5 IoU@0.3 Acc ANLS GQA@0.5 GQA@0.3
  w/o Trans. Layer 16.36 25.72 9.18 0.1272 2.14 2.42
w/ 1-layer Trans. 25.08 38.18 8.70 0.1262 2.87 4.23
w/ 2-layer Trans. 28.60 41.94 10.45 0.1520 3.50 4.23
w/ 3-layer Trans. 27.00 39.53 8.97 0.1294 3.17 3.89
@@ V-C Ablation Study (RQ2)
Effectiveness of Temporal-to-Spatial Grounding. In Tab. V, we respectively study the effectiveness of T2S-QA by removing temporal and spatial grounding (“w/o TSG”), only removing temporal grounding (“w/o TG”), and only removing spatial grounding (“w/o SG”). Specifically, “w/o TG” variant uses the grounded OCR tokens in all video frames as positive OCR tokens and the others as negative OCR tokens; “w/o SG” variant uses all OCR tokens of the positive video frame as the positive OCR token, and all OCR tokens of the negative video frame as the negative OCR token; “w/o TSG” variant represents reserving all video frames and OCR tokens for answering, i.e., without answer grounding.
Tab. V shows that both spatial and temporal grounding are key for better QA performance. We note that the model variant “w/o SG” performs the worst in both answer grounding and grounded QA. This suggests that only learning the differences between video frames is insufficient to capture subtle local changes in scene text within a frame, making it challenging to achieve accurate answer grounding. Moreover, compared with T2S-QA, “w/o TG” also shows a significant decrease of 13.29% on IoU@0.5 and 1% decrease in GQA@0.5. This indicates that directly performing one-stage grounding of scene text across the entire video yields suboptimal results.
Study Inputs of Answer Decoder. Tab. VI shows that when removing OCR tokens (“w/o O”) both T2S-QA’s grounding and QA results degenerate significantly by 11.16% on IoU@0.5 and 4.76% in Acc respectively. This demonstrates the significance of OCR token grounding for answer prediction. Additionally, we show that both the frame appearance F feature and question Q help the performance.
@@
Study of Scene Text Features. In Tab. VII, the results show that linguistic features obtained by character-level encoding are crucial based on the scene text recognition results. Capturing spatial and temporal features can complement the spatiotemporal characteristics of scene text in videos. Their combination produces the best performance.
Different Numbers of Video Frames. We conduct experiments with four different numbers of sampled video frames, i.e., $T$ = 32, 48, 64, and 80. Tab. VIII shows that uniformly sampling 80 frames can cover over 96% of the annotated scene-text regions. However, the best results are achieved at frame number of 64. We speculate that learning with too many frames make it harder to identify the very limited positive frames that carry the needed scene texts.
Different Numbers of Transformer Layers. In Tab. IX, we conduct experiments by varying the number of Transformer layers for the architecture introduced at Sec. III-A. It shows that when the Transformer module is removed (“w/o Trans. Layer”), the performance on IoU@0.5 drops by 12.24%, and the performance on GQA@0.5 drops by 1.03%. The results show that the optimal configuration is with two layers.
TABLE X: Study of scene-text grounding and recognition in Grounded TextVideoQA.
@@
TABLE XI: Performance of T2S-QA under different scene-text detection and recognition systems.
  OCR Detection OCR Recognition IoU@0.5 Acc ANLS GQA@0.5
  GoMatching [19] ABINet [12] 25.26 9.89 0.1436 2.15
TransVTSpotter [46] ABINet [12] 28.14 10.20 0.1472 2.69
TransVTSpotter [46] CLIPOCR [45] 28.60 10.45 0.1520 3.50
@@ V-D In-Depth Analysis (RQ3)
Our previous experiments show that models obtain low QA accuracy even with relatively higher grounding behavior. We speculate that the reasons are two folds: ineffective answer evaluation and poor scene-text detection and recognition. The former indicates that the predicted textual answer does not exactly match the ground-truth one (e.g. 30 vs. 30 M.P.H in Fig. 1). The latter hints that the OCR system fails to convert the grounded scene-text into textual answers (e.g. M vs. M.P.H).
@@
Human Performance and Upper Bound. To discern the error from ineffective answer evaluation, we sample around 20% test data which contains 200 questions and 163 videos and conduct a human study by asking volunteers to answer the questions and ground the related scene text areas. Table X shows that even human (we assume that humans are with perfect scene-text recognition) can only achieve an accuracy of 64.32%. A further investigation of the the wrong answers suggests that an absolute 25% out of 36% of they are acceptable and should be regarded as correct. For example, “35131.86” vs. “35,131.86”. The results suggest that stringent string matching is ineffective in assessing TextVQA models. Yet, human’s significantly higher grounding performance indicates that grounding annotation is less subjective, highlighting the importance of directly evaluating scene-text regions. The gap between our model and human performance underscores the need for further research into trustworthy TextVideoQA.
@@
We report the “Upper Bound” result with the TransVTSpotter [46] detection and CLIPOCR [45] recognition systems, obtained via brute-force search over all detected text regions and recognized outputs. The result in Table X suggests that even a brute-force search can only achieve a grounding accuracy of 56.52%. Note that T2S-QA achieves 28.60%. This reflects the effectiveness of our grounding approach, though the accuracy is not so high. Additionally, we find that the “Upper Bound” is inferior to human in grounding but exceeds human in QA accuracy. This reveals the inadequacy of the existing OCR system in scene-text detection and the challenge of effectively evaluating textual answers because of human subjectivity.
Impact of Scene Text Detection and Recognition. As shown in Table XI, we evaluate various scene text detection and recognition systems and analyze their impact on the Grounded TextVideoQA task from three perspectives. 1) To assess the impact of OCR detection systems, we test the variant where the ground-truth scene text for each QA pair is directly provided to the answer decoder. Compared to the human-annotated answer grounding result (“w/ CLIPOCR GT Box”), the performance of directly iterating over all OCR detection results (“Upper Bound”) drops by 40.18% on IoU@0.5. This drop illustrates the difficulty of scene text detection in dynamic video environments. 2) We analyze the impact of OCR recognition systems. The GQA@0.5 score of “T2S-QA (CLIPOCR)” is 25.1% which is lower than its IoU@0.5 score, suggesting incorrect textual answers even when the scene text is located accurately. The comparison between “T2S-QA (CLIPOCR)” and “w/ CLIPOCR GT Box” shows only marginal improvement (1.51% in Acc) despite perfect grounding. These results emphasize the limitations of current video-based scene text recognition systems. 3) We evaluate various combinations of OCR detection and recognition systems. Using the same recognition system, TransVTSpotter achieves superior grounding due to its robust detection capabilities. Similarly, employing CLIPOCR as the recognition method significantly improves grounding accuracy when using a fixed detection system. These results demonstrate that high-quality detection-recognition combinations enhance grounding performance and optimize QA accuracy.
Figure 7: Case-Study of T2S-QA on ViTXT-GQA test set. The green and yellow boxes in video frames denote the ground-truth boxes and the model-grounded OCR tokens. The blue bar denotes the annotated temporal segment.